OpenStudy (anonymous):

What is the slope-intercept form for (-4,1) and (2,4)?

OpenStudy (anonymous):

Find the slope, \[m=\frac{ y _{2} - y _{1}}{ x _{2} - x _{1}}\] \[m=\frac{ 4 - 1 }{ 2+4 }\] \[m=\frac{ 5 }{ 6 }\] Now I have the slope and two points. I know I can find the equation (by solving first for "b") if I have a point and the slope. So I need to pick one of the points (it doesn't matter which one), and use it to solve for b. Using the point (-4,1), I get: y = mx + b 1 = (5/6)(–4) + b 1 = 20/6 + b 1 – 20/6 = b 6/6 – 20/6 = b b = 14/6 ...so y = ( 5/6 ) x + 14/6. You should get the same answer no matter which point you choose \[y=\frac{ 5 }{ 6 }x+\frac{ 14 }{ 6 }\]

OpenStudy (anonymous):

On the other hand, if I use the point (2,4), I get: y = mx + b 4 = (5/6)(2) + b 4 = 10/6 + b 4 + 10/6 = b 24/6 - 10/6 = b b = 14/6

OpenStudy (dean.shyy):

This may help: http://is.gd/B9KWed